write an inequality and solve for the homework average.
Claire's test average is 90.The test average is 2/3 of her final grade, and her homework average is 1/3 of her final grade.What homework average does she need in order to have a final grade of at least 93?

Answer by checkley71(8403)

(2/3)90+(1/3)X>=93
60+X/3>=93
X/3>=93-60
X/3>=33
X>=33/3
X>=100 THIS SAYS SHE NEED A PERFECT SCORE ON HER HOMEWORK.